VEGAS Pulsar Project Meeting: 2015 September 21st 2:00 - 3:00pm ET

Room / Connection Details

  • GB-137 / Soc 280
  • 192.33.117.12##7144
  • 434-817-6443

Housekeeping Details

  • Main wiki page
  • gbsapp Mailing List
  • Location and regular date and time of meeting - agreed weekly on Mondays; GB-137 / Soc-280; 2-3pm ET
  • ETK codes:
    • Electronics: 432510.GB3015
    • Software: 893252.GB3015
    • Science: 432540.GB3015

Resource Allocations

GUPITA, ARINDAM 1.00
BRANDT, JOE 0.1
FORD, JOHN 0.05
MCCULLOUGH, RANDY 0.05
RAY, JASON 0.55
BLOSS, MARTY 0.1
CREAGER, RAY 0.65
RICHMOND DECKER, JUSTIN 0.2
LYNCH, RYAN 0.25
DEMOREST, PAUL 0.1
PRESTAGE, RICHARD 0.05

Present

  • Randy, Ray, Justin, Richard, John, Jason, Vereese, Ryan, Paul

Agenda

General Updates

  • BOF Files
    • New c1500x0512 mode is now available after errors found in a previous version
      • Ryan will update vegas.conf and BOF file status page with new file names and test
    • Randy still struggling with timing errors in c1500x4096 mode but is making progress
  • Software
    • We found an error in the way fftshift was being set because the vegas.conf file was being read from the local directory when present instead of the etc/config directory where the official copy is kept. This has now been fixed.
    • Ryan and Justin will be testing changes to the manager for coherent mode tomorrow
    • Ray still working on spectral line modes for Matrix
  • Wideband AP
    • Circuit board has been designed and will be turned over for cam file generation this week
    • Vereese will work on controller code and diagnostics once hardware is available

Discussion

  • Richard would like to compile a list of updates/new info for his visit to SHAO
    • Rebuild BOF files under the latest version of Xilinx and CASPER libraries
      • This will require minor changes to the dealer/player to handle new register naming conventions
    • Updates to guppi_daq to fix polyco generation in cal mode (This is currently up-to-date in Paul's github repo)
    • Builds with adc snap blocks to use for power monitor
    • Adding support for switching signal generation
    • Joe made updates to net-task unpacking to enable wide BW spectral line modes
      • Does this affect pulsar modes
    • Low bandwidth BOFs when available
  • Quarterly planning planning
    • Highest priority is finishing c1500x4096 mode and making sure that BL has a basic software interface for taking data
    • We will request more of Jason and Randy's time, with Randy's allocation from VEGAS spectral line converted to VPM
    • Justin to work on config tool
      • Allocation will remain the same
      • We will request 0.05 of Melinda so that she is available to answer questions
    • Vereese will need some time to work on wideband AP but most of that work will likely happen next quarter. Some development can happen now, and she can help with floor planning. Her highest priority is PAF but we will request 0.1 of her time so she can start coming up to speed.

Date / Time of Next Meeting

  • September 28
Topic revision: r1 - 2015-10-12, RyanLynch
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback